OpenStudy (soahelp):

MEDAL!!!!! Two equations are given below: a − 3b = 4 a = b − 2 What is the solution to the set of equations in the form (a, b)? (−2, −2) (−3, −1) (−9, −7) (−5, −3)

OpenStudy (soahelp):

EXPLAIN your answer to plz

OpenStudy (welshfella):

a - 3b = 4 a - b = -2 if you subtract these you eliminate a and get an equation in b only

OpenStudy (welshfella):

can you do that /

OpenStudy (soahelp):

how do i subtract it

OpenStudy (welshfella):

a - a = 0 of course -3b - (-b) = ? 4 - (-2) = ?

OpenStudy (babynini):

@SoaHelp Try putting the points into your equation and see which comes out true :)

OpenStudy (welshfella):

can you work these out?

OpenStudy (soahelp):

0B 2

OpenStudy (welshfella):

you can do it Babyninis way but that won't help you to learn how to solve these type of questions

OpenStudy (babynini):

a − 3b = 4 a = b − 2 What is the solution to the set of equations in the form (a, b)? a) (-2)-3(-1)=4 true or false? b) (−3)-3(−1)=4 true or false? and so on until you find your correct answer! When you find it, plug in the points to the second equation just to double check :)

OpenStudy (welshfella):

-3b - (-b) = -3b + b (the consecutive - sign is same as a +

OpenStudy (babynini):

noo, not the diamonds! :|

OpenStudy (soahelp):

im so confused

OpenStudy (welshfella):

what is -3b + b ?

OpenStudy (soahelp):

3

OpenStudy (welshfella):

no you;ve lost the b! 2b + b = 3b -3b + b = -2b

OpenStudy (soahelp):

1

OpenStudy (soahelp):

1b

OpenStudy (welshfella):